OpenStudy (anonymous):

Which statement describes the translation of the graph of y = −2(x − 3)2 + 5 from standard position? a)Moved up and to the right b)Moved up and to the left c)Moved down and to the right d)Moved down and to the left

OpenStudy (anonymous):

@Noland__DeWitte see if you can figure this one out :)?

OpenStudy (anonymous):

brb k?

OpenStudy (anonymous):

Alrighty

OpenStudy (mathmate):

If the standard graph is \(f(x)=x^2\), then the modified graph is \(g(x)=a(x-h)^2+k\) where a=vertical scaling factor, stretches when a>1 compresses when 0<a<1 stretches and reflects about the x-axis when a<-1 compresses when -1<a<0 h=horizontal translation (positive to the right) k=vertical translation (positive upwards) Example: Find translation of \(y=3(x+5)^2-2\) First rewrite it as \(y=3(x-(-5))^2-2\) which gives a=3, h=-5, k=-2 so the graph of \(y=x^2\) has been stretched vertically by a factor of 3, translated to the left by 5 (h=-5) , and downwards by 2 (k=-2). The results of the \(example\) are shown in the following attachment.

OpenStudy (anonymous):

ahh ok..so would the answer to my original question be d?

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Parenthesis lie when translating graphs, negatives become positives and positives become negatives. Which means, -3 would be +3 and +5 would stay the same. Therefore, the graph would move up and to the right.

OpenStudy (anonymous):

I hope you're right haha thanks

OpenStudy (anonymous):

I'm 99% sure

OpenStudy (anonymous):

could you help me with another one like this?

OpenStudy (mathmate):

@gypsy1234 If you would rewrite the expression in the form \(y=a(x-h)^2+k\) you will find that h=3, k=5, a=-2, so that in addition to flipping upside-down, stretched vertically, the vertex is translated by (3,5), and everything will make sense. I would put @FireWolfSpirit 's statement "Therefore, the \(graph\) would move up and to the right." more as "Therefore, the \(vertex\) would move up and to the right." since the graph is flipped.